# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 13. Appendices
13.1. Topology of the unlabeled and labeled path spaces.
@@
Let $\mathsf{S}_{\mathrm{s,i}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\mathbb{R}^{d})=\infty,\mathsf{s}(\{s\})\in\{0,1\}\text{ for all }s\in\mathbb{R}^{d}\}$. Let $W(\mathsf{S}_{\mathrm{s,i}})$ be the set of all $\mathsf{S}_{\mathrm{s,i}}$-valued continuous paths defined on $0,\infty)$. We quote:
Lemma 13.1 ([28, Lem. 7.1]).
Each $\mathsf{w}=\{\mathsf{w}_{t}\}\in W(\mathsf{S}_{\mathrm{s,i}})$ can be written as
$\mathsf{w}_{t}=\sum_{i=1}^{\infty}\delta_{w^{i}(t)},\quad w^{i}=\{w^{i}(t)\}\in C(I_{i};\mathbb{R}^{d}).$
Here, $I_{i}$ is an interval of the form $0,b_{i})$ or $(a_{i},b_{i})$, where $0\leq a_{i}<b_{i}\leq\infty$. For a given $\mathsf{w}\in W(\mathsf{S}_{\mathrm{s,i}})$, the set $\{(w^{i},I_{i}),i\in\mathbb{N}\}$ is uniquely determined (except labeling).
@@
For $\mathbf{w}=(w^{i})_{i=1}^{\infty}$, we set $\mathfrak{u}_{\mathrm{path}}(\mathbf{w})=\mathsf{w}$, where $\mathsf{w}_{t}=\sum_{i=1}^{\infty}\delta_{w^{i}(t)}$.
Lemma 13.2 ([28, Lem. 7.2]).
@@
For a label $\mathfrak{l}$, we have a unique $\mathfrak{l}_{\mathrm{path}}\!:\!W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})\!\to\!C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$ such that $\mathfrak{l}(\mathsf{w}_{0})=\mathfrak{l}_{\mathrm{path}}(\mathsf{w})_{0}$ and $\mathfrak{u}_{\mathrm{path}}\circ\mathfrak{l}_{\mathrm{path}}(\mathsf{w})=\mathsf{w}$.
Let $\|u-v\|_{T}=\sup_{0\leq t\leq T}|u(t)-v(t)|$, $\mathbf{u}=(u^{i})_{i=1}^{\infty}$, and $\mathbf{v}=(v^{i})_{i=1}^{\infty}$. We set
@@
$\varrho_{\mathrm{path}}(u,v)=\sum_{T=1}^{\infty}\frac{1}{2^{T}}(1\wedge\|u-v\|_{T}),\quad\varrho_{\mathrm{lpath}}(\mathbf{u},\mathbf{v})=\sum_{i=1}^{\infty}\frac{1}{2^{i}}(1\wedge\varrho_{\mathrm{path}}(u^{i},v^{i})).$
We endow $C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$ with $\varrho_{\mathrm{lpath}}$. $C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$ is a complete separable metric space under $\varrho_{\mathrm{lpath}}$. Let $\varrho_{\mathrm{Proho}}$ be the Proholov metric on $\mathsf{S}$:
$\varrho_{\mathrm{Proho}}(\mathsf{s},\mathsf{s}^{\prime})=\inf\{\epsilon>0;\mathsf{s}(B)\leq\mathsf{s}^{\prime}(B^{\epsilon})+\epsilon,\mathsf{s}^{\prime}(B)\leq\mathsf{s}(B^{\epsilon})+\epsilon\}.$
Here, $B^{\epsilon}$ is the open $\epsilon$-neighborhood of $B$. Let $f_{R}\in C_{0}(\mathbb{R}^{d})$ be such that $0\leq f_{R}\leq 1$, $f_{R}(x)=1$ on $S_{R}$, and $f_{R+1}(x)=0$ on $S_{R+1}^{c}$. Let $f\mathsf{s}=\sum_{i}f(s^{i})\delta_{s^{i}}$ be the measure $f\mathsf{s}$ on $\mathbb{R}^{d}$. Then, the vague topology on $\mathsf{S}$ is given by the separable, complete metric defined by
$\varrho_{\mathrm{vague}}(\mathsf{s},\mathsf{s}^{\prime})=\sum_{R=1}^{\infty}\frac{1}{2^{R}}\Big{(}1\wedge\varrho_{\mathrm{Proho}}(f_{R}\mathsf{s},f_{R}\mathsf{s}^{\prime})\Big{)}.$
We refer to [12] for these metrics. We set
$\varrho_{\mathrm{upath}}(\mathsf{w},\mathsf{w}^{\prime})=\sum_{T=1}^{\infty}\frac{1}{2^{T}}\Big{(}1\wedge\sup_{0\leq t\leq T}\varrho_{\mathrm{vague}}(\mathsf{w}_{t},\mathsf{w}_{t}^{\prime})\Big{)}.$
@@
Then, $C([0,\infty);\mathsf{S})$ is a complete separable metric space under $\varrho_{\mathrm{upath}}$. We set
$\text{$\mathcal{NBJ}=\{\mathbf{w}\in C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}};\text{$I_{R,T}(\mathbf{w})<\infty$ for each $R,T\in\mathbb{N}$}\}$},$
where $I_{R,T}$ is defined by (2.54).
Lemma 13.3 ([28, Lem. 7.3]).
@@
13.2. Closability and extension of Dirichlet forms and strongly local, quasi-regular Dirichlet forms.
In Subsection 13.2, we quote the concepts of closability and extension of Dirichlet forms, strongly local, quasi-regular Dirichlet forms, and related quantities following [5, 7]. We repeatedly use them in the present paper.
@@
For non-negative symmetric bilinear forms $(\mathscr{E}_{1},\mathscr{D}_{1})$ and $(\mathscr{E}_{2},\mathscr{D}_{2})$ on $L^{2}(\nu)$, we say $(\mathscr{E}_{2},\mathscr{D}_{2})$ is an extension of $(\mathscr{E}_{1},\mathscr{D}_{1})$ if
@@
We write $(\mathscr{E}_{1},\mathscr{D}_{1})\prec(\mathscr{E}_{2},\mathscr{D}_{2})$ if $(\mathscr{E}_{2},\mathscr{D}_{2})$ is an extension of $(\mathscr{E}_{1},\mathscr{D}_{1})$ on $L^{2}(\nu)$.
@@
Definition 13.1.
[Closable and closure [5, 7]] i A non-negative symmetric bilinear form $(\mathscr{E},\mathscr{D}_{0})$ densely defined on $L^{2}(\nu)$ is called closable on $L^{2}(\nu)$ if, for any $\mathscr{E}$-Cauchy sequence $f_{n}\in\mathscr{D}_{0}$ such that $\lim\|f_{n}\|_{L^{2}(\nu)}=0$ as $n\to\infty$, it holds that $\lim_{n\to\infty}\mathscr{E}(f_{n},f_{n})=0$.
ii If $(\mathscr{E},\mathscr{D}_{0})$ is closable on $L^{2}(\nu)$, then there exists a closed extension of $(\mathscr{E},\mathscr{D}_{0})$. The smallest closed extension $(\mathscr{E},\mathscr{D})$ of $(\mathscr{E},\mathscr{D}_{0})$ is called the closure of $(\mathscr{E},\mathscr{D}_{0})$.
The following simple fact is used repeatedly in the present paper.
@@
Let $(\mathscr{E}_{1},\mathscr{D}_{1})$ and $(\mathscr{E}_{2},\mathscr{D}_{2})$ be non-negative symmetric bilinear forms on $L^{2}(\nu)$. Let $(\mathscr{E}_{2},\mathscr{D}_{2})$ be an extension of $(\mathscr{E}_{1},\mathscr{D}_{1})$. Let $(\mathscr{E}_{2},\mathscr{D}_{2})$ be closable on $L^{2}(\nu)$. Then, $(\mathscr{E}_{1},\mathscr{D}_{1})$ is closable on $L^{2}(\nu)$.
@@
If $\{f_{n}\}$ is a Cauchy sequence of $(\mathscr{E}_{1},\mathscr{D}_{1})$, then $\{f_{n}\}$ is a Cauchy sequence of $(\mathscr{E}_{2},\mathscr{D}_{2})$ by (13.1). Then, $\lim_{n\to\infty}\mathscr{E}_{2}(f_{n},f_{n})=0$ because of the closability of $(\mathscr{E}_{2},\mathscr{D}_{2})$ on $L^{2}(\nu)$. Hence, $\lim_{n\to\infty}\mathscr{E}_{1}(f_{n},f_{n})=0$ from (13.1). This completes the proof. ∎
A sequence of closed non-negative bilinear forms $(\mathscr{E}^{n},\mathscr{D}^{n})$ on $L^{2}(\nu)$ is said to converge to $(\mathscr{E},\mathscr{D})$ on $L^{2}(\nu)$ in the strong resolvent sense if their $\alpha$-resolvents $G_{\alpha}^{n}$ converge strongly to the $\alpha$-resolvent $G_{\alpha}$ of $(\mathscr{E},\mathscr{D})$ in $L^{2}(\nu)$ for each $\alpha>0$ (cf. [38]). This implies that the $L^{2}(\nu)$-semi-groups $T_{t}^{n}$ associated with $(\mathscr{E}^{n},\mathscr{D}^{n})$ on $L^{2}(\nu)$ converge strongly to the $L^{2}(\nu)$-semigroup $T_{t}$ associated with $(\mathscr{E},\mathscr{D})$ on $L^{2}(\nu)$ for each $t\geq 0$.
@@
$\mathscr{E}^{\infty}(f,f)=\lim_{n\to\infty}\mathscr{E}^{n}(f,f),\quad\mathscr{D}^{\infty}=\{f\in\bigcap_{n\in\mathbb{N}}\mathscr{D}^{n}\,;\,\lim_{n\to\infty}\mathscr{E}^{n}(f,f)<\infty\}.$ (13.2)
Then, $(\mathscr{E}^{\infty},\mathscr{D}^{\infty})$ is a closed symmetric bilinear form on $L^{2}(\nu)$. $\{(\mathscr{E}^{n},\mathscr{D}^{n})\}_{n\in\mathbb{N}}$ converges to $(\mathscr{E}^{\infty},\mathscr{D}^{\infty})$ in the strong resolvent sense on $L^{2}(\nu)$.
@@
Then, $\{(\mathscr{E}^{n},\mathscr{D}^{n})\}_{n\in\mathbb{N}}$ converges to $(\mathscr{E}_{\mathrm{max}}^{\infty},\mathscr{D}_{\mathrm{max}}^{\infty})$ in the strong resolvent sense on $L^{2}(\nu)$. Here, $(\mathscr{E}_{\mathrm{max}}^{\infty},\mathscr{D}_{\mathrm{max}}^{\infty})$ is the maximal closable part of $(\mathscr{E}_{\mathrm{max}}^{\infty},\mathscr{D}_{\mathrm{max}}^{\infty})$ on $L^{2}(\nu)$.
@@
We say a closed non-negative symmetric bilinear form $(\mathscr{E},\mathscr{D})$ on $L^{2}(\nu)$ is a symmetric Dirichlet form if any $u\in\mathscr{D}$ satisfies
@@
Let $\mathcal{O}$ be the family of all open subsets of $S$. For ${A}\in\mathcal{O}$, let $\mathcal{L}_{{A},1}=\{f\in\mathscr{D};f\geq 1\ \mu\text{-a.e.\,on }{A}\}$. We set $\mathcal{O}_{0}=\{{A}\in\mathcal{O};\mathcal{L}_{{A},1}\neq\emptyset\}$. Let $\mathscr{E}_{1}=\mathscr{E}+(\cdot,*)_{L^{2}(\nu)}$. For an open set ${A}\in\mathcal{O}$, we set
@@
An increasing sequence of closed sets $\{F_{k}\}$ is called an $\mathscr{E}$-nest if $\cup_{k\geq 1}\mathscr{D}(F_{k})$ is $\mathscr{E}_{1}$-dense, where $\mathscr{D}(F_{k})=\{f\in\mathscr{D};f=0\text{ $\nu$-a.e.\,on }F_{k}^{c}\}$ and $\mathscr{E}_{1}=\mathscr{E}+(\cdot,\cdot)_{L^{2}(\nu)}$. A function $f$ is called $\mathscr{E}$-quasi-continuous if for any $\epsilon>0$, there exists an open set $\mathcal{O}$ with $\mathrm{Cap}(\mathcal{O})<\epsilon$ such that $f|_{S\backslash\mathcal{O}}$ is finite and continuous. We call a subset $\mathfrak{N}\subset S$ an $\mathscr{E}$-polar set if there exists an $\mathscr{E}$-nest $\{F_{k}\}$ such that $\mathfrak{N}\subset\cap_{k}(S\backslash F_{k})$.
@@
A Dirichlet form $(\mathscr{E},\mathscr{D})$ on $L^{2}(\nu)$ is called quasi-regular if:
((QR1)) there exists an $\mathscr{E}$-nest $\{F_{k},k\geq 1\}$ consisting of compact sets;
((QR2)) there exists $\mathscr{E}+(\cdot,*)_{L^{2}(\nu)}$-dense subset of $\mathscr{D}$ whose elements have $\mathscr{E}$-quasi-continuous $\mu$-version;
((QR3)) there exists $\{f_{k},k\geq 1\}\subset\mathscr{D}$ having $\mathscr{E}$-quasi-continuous $\mu$-versions $\{\tilde{f}_{k},k\geq 1\}\subset\mathscr{D}$ and an $\mathscr{E}$-polar set $\mathfrak{N}\subset S$ such that $\{\tilde{f}_{k},k\geq 1\}\subset\mathscr{D}$ separates the points on $S\backslash\mathfrak{N}$.

### v4

@@ 13. Appendices
13.1. Topology of the unlabeled and labeled path spaces
@@
Let $\mathsf{S}_{\mathrm{s,i}}=\{\mathsf{s}\in\mathsf{S}\,;\,\mathsf{s}(\mathbb{R}^{d})=\infty,\mathsf{s}(\{s\})\in\{0,1\}\text{ for all }s\in\mathbb{R}^{d}\}$. Let $W(\mathsf{S}_{\mathrm{s,i}})$ be the set of all $\mathsf{S}_{\mathrm{s,i}}$-valued continuous paths defined on $0,\infty)$.
Remark 13.1.
The preprint [30] is a longer version of the published paper [31], containing additional topological and labeling results that were omitted from the journal version.
Lemma 13.1 ([30, Lem. 7.1]).
Let $I_{i}$ be of the form $0,b_{i})$ or $(a_{i},b_{i})$ with $0\leq a_{i}<b_{i}\leq\infty$. Each $\mathsf{w}\in W(\mathsf{S}_{\mathrm{s,i}})$ admits a representation
$\mathsf{w}_{t}=\sum_{i=1}^{\infty}\delta_{w^{i}(t)},\quad w^{i}\in C(I_{i};\mathbb{R}^{d}).$
The collection $\{(w^{i},I_{i})\}_{i\in\mathbb{N}}$ is uniquely determined up to relabeling.
@@
For $\mathbf{w}=(w^{i})_{i=1}^{\infty}$, define $\mathfrak{u}_{\mathrm{path}}(\mathbf{w})=\mathsf{w}$ by $\mathsf{w}_{t}=\sum_{i=1}^{\infty}\delta_{w^{i}(t)}$.
Lemma 13.2 ([30, Lem. 7.2]).
@@
For any label $\mathfrak{l}$, there exists a unique map $\mathfrak{l}_{\mathrm{path}}:W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})\to W(\mathbb{R}^{d})^{\mathbb{N}}$ such that $\mathfrak{l}(\mathsf{w}_{0})=\mathfrak{l}_{\mathrm{path}}(\mathsf{w})_{0}$ and $\mathfrak{u}_{\mathrm{path}}\circ\mathfrak{l}_{\mathrm{path}}(\mathsf{w})=\mathsf{w}$.
Set $\|u\|_{T}=\sup_{0\leq t\leq T}|u(t)|$. For $\mathbf{u}=(u^{i})_{i\in\mathbb{N}}$ and $\mathbf{v}=(v^{i})_{i\in\mathbb{N}}$, define
@@
$\varrho_{\mathrm{path}}(u,v)=\sum_{T=1}^{\infty}\frac{1}{2^{T}}(1\wedge\|u-v\|_{T}),\quad\varrho_{\mathrm{lpath}}(\mathbf{u},\mathbf{v})=\sum_{i=1}^{\infty}\frac{1}{2^{i}}\bigl(1\wedge\varrho_{\mathrm{path}}(u^{i},v^{i})\bigr).$
@@
Then $(W(\mathbb{R}^{d})^{\mathbb{N}},\varrho_{\mathrm{lpath}})$ is a complete separable metric space.
Let $f_{R}\in C_{0}(\mathbb{R}^{d})$ satisfy $0\leq f_{R}\leq 1$, $f_{R}=1$ on $S_{R}$, and $f_{R+1}=0$ on $S_{R+1}^{c}$. For $\mathsf{s}=\sum_{i}\delta_{s^{i}}$, set $f_{R}\mathsf{s}=\sum_{i}f_{R}(s^{i})\delta_{s^{i}}$. Let $\varrho_{\mathrm{Prh}}$ denote the Prohorov metric on $\mathsf{S}$ [11, p. 11]. The vague topology on $\mathsf{S}$ is induced by the complete separable metric
$\varrho_{\mathrm{vague}}(\mathsf{s},\mathsf{s}^{\prime})=\sum_{R=1}^{\infty}\frac{1}{2^{R}}\bigl(1\wedge\varrho_{\mathrm{Prh}}(f_{R}\mathsf{s},f_{R}\mathsf{s}^{\prime})\bigr).$
Then $C([0,\infty);\mathsf{S})$ is a complete separable metric space under the metric
$\varrho_{\mathrm{upath}}(\mathsf{w},\mathsf{w}^{\prime})=\sum_{T=1}^{\infty}\frac{1}{2^{T}}\bigl(1\wedge\sup_{0\leq t\leq T}\varrho_{\mathrm{vague}}(\mathsf{w}_{t},\mathsf{w}_{t}^{\prime})\bigr).$
Let $I_{R,T}$ be defined in (2.40), and set
$\mathcal{NBJ}=\bigl\{\mathbf{w}\in W(\mathbb{R}^{d})^{\mathbb{N}};\ I_{R,T}(\mathbf{w})<\infty\ \text{for all }R,T\in\mathbb{N}\bigr\}.$
Lemma 13.3 ([30, Lem. 7.3]).
@@
13.2. Closability and extensions of Dirichlet forms, and strongly local, quasi-regular Dirichlet Forms
In this subsection, we recall notions related to Dirichlet forms from [5, 8].
@@ 13.2. Closability and extensions of Dirichlet forms, and strongly local, quasi-regular Dirichlet Forms
For non-negative symmetric bilinear forms $(\mathscr{E}_{1},\mathscr{D}_{1})$ and $(\mathscr{E}_{2},\mathscr{D}_{2})$, we say $(\mathscr{E}_{2},\mathscr{D}_{2})$ is an extension of $(\mathscr{E}_{1},\mathscr{D}_{1})$ if
@@
Definition 13.1 (Closable and closure [8]).
i A non-negative symmetric bilinear form $(\mathscr{E},\mathscr{D}_{0})$ densely defined on $L^{2}(\nu)$ is called closable on $L^{2}(\nu)$ if, for any $\mathscr{E}$-Cauchy sequence $f_{n}\in\mathscr{D}_{0}$ such that $\lim\|f_{n}\|_{L^{2}(\nu)}=0$ as $n\to\infty$, it holds that $\lim_{n\to\infty}\mathscr{E}(f_{n},f_{n})=0$.
ii If $(\mathscr{E},\mathscr{D}_{0})$ is closable on $L^{2}(\nu)$, then there exists a closed extension of $(\mathscr{E},\mathscr{D}_{0})$. The smallest closed extension $(\mathscr{E},\mathscr{D})$ is called the closure of $(\mathscr{E},\mathscr{D}_{0})$.
@@
Let $(\mathscr{E}_{1},\mathscr{D}_{1})$ and $(\mathscr{E}_{2},\mathscr{D}_{2})$ be non-negative symmetric bilinear forms on $L^{2}(\nu)$. Let $(\mathscr{E}_{2},\mathscr{D}_{2})$ be an extension of $(\mathscr{E}_{1},\mathscr{D}_{1})$. Let $(\mathscr{E}_{2},\mathscr{D}_{2})$ be closable on $L^{2}(\nu)$. It holds that $(\mathscr{E}_{1},\mathscr{D}_{1})$ is closable on $L^{2}(\nu)$.
@@
If $\{f_{n}\}$ is a Cauchy sequence of $(\mathscr{E}_{1},\mathscr{D}_{1})$, then $\{f_{n}\}$ is a Cauchy sequence of $(\mathscr{E}_{2},\mathscr{D}_{2})$ by (13.1). Then $\lim_{n\to\infty}\mathscr{E}_{2}(f_{n},f_{n})=0$ because of the closability of $(\mathscr{E}_{2},\mathscr{D}_{2})$ on $L^{2}(\nu)$. Hence, $\lim_{n\to\infty}\mathscr{E}_{1}(f_{n},f_{n})=0$ from (13.1). This completes the proof. ∎
A sequence of closed, nonnegative bilinear forms $(\mathscr{E}^{n},\mathscr{D}^{n})$ on $L^{2}(\nu)$ is said to converge to $(\mathscr{E},\mathscr{D})$ on $L^{2}(\nu)$ in the strong resolvent sense if the associated $\alpha$-resolvents $G_{\alpha}^{n}$ converge strongly in $L^{2}(\nu)$ to the $\alpha$-resolvent $G_{\alpha}$ of $(\mathscr{E},\mathscr{D})$ for each $\alpha>0$. This is equivalent to the same convergence holding for some $\alpha>0$ (cf. [37]).
In this case, the $L^{2}(\nu)$-semigroups $T_{t}^{n}$ associated with $(\mathscr{E}^{n},\mathscr{D}^{n})$ converge strongly in $L^{2}(\nu)$ to the $L^{2}(\nu)$-semigroup $T_{t}$ associated with $(\mathscr{E},\mathscr{D})$ for each $t\geq 0$.
@@
$\mathscr{E}^{\infty}(f,f)=\lim_{n\to\infty}\mathscr{E}^{n}(f,f),\quad\mathscr{D}^{\infty}=\{f\in\bigcap_{n\in\mathbb{N}}\mathscr{D}^{n}\,;\,\lim_{n\to\infty}\mathscr{E}^{n}(f,f)<\infty\}.$
It holds that $(\mathscr{E}^{\infty},\mathscr{D}^{\infty})$ is a closed symmetric bilinear form on $L^{2}(\nu)$. $\{(\mathscr{E}^{n},\mathscr{D}^{n})\}_{n\in\mathbb{N}}$ converges to $(\mathscr{E}^{\infty},\mathscr{D}^{\infty})$ in the strong resolvent sense on $L^{2}(\nu)$.
@@
It holds that $\{(\mathscr{E}^{n},\mathscr{D}^{n})\}_{n\in\mathbb{N}}$ converges to $(\mathscr{E}_{\mathrm{max}}^{\infty},\mathscr{D}_{\mathrm{max}}^{\infty})$ in the strong resolvent sense on $L^{2}(\nu)$. Here, $(\mathscr{E}_{\mathrm{max}}^{\infty},\mathscr{D}_{\mathrm{max}}^{\infty})$ is the maximal closable part of $(\mathscr{E}_{\mathrm{max}}^{\infty},\mathscr{D}_{\mathrm{max}}^{\infty})$ on $L^{2}(\nu)$.
@@
We say a closed non-negative symmetric bilinear form $(\mathscr{E},\mathscr{D})$ densely defined on $L^{2}(\nu)$ is a symmetric Dirichlet form if any $u\in\mathscr{D}$ satisfies
@@
Let $\mathcal{O}$ be the family of all open subsets of $S$. For ${A}\in\mathcal{O}$, let $\mathcal{L}_{{A},1}=\{f\in\mathscr{D};f\geq 1\ \nu\text{-a.e.\,on }{A}\}$. We set $\mathcal{O}_{0}=\{{A}\in\mathcal{O};\mathcal{L}_{{A},1}\neq\emptyset\}$. Let $\mathscr{E}_{1}=\mathscr{E}+(\cdot,*)_{L^{2}(\nu)}$. For an open set ${A}\in\mathcal{O}$, we set
@@
Let $\mathscr{D}(F_{k})=\{f\in\mathscr{D};f=0\text{ $\nu$-a.e.\,on }F_{k}^{c}\}$ and $\mathscr{E}_{1}=\mathscr{E}+(\cdot,\cdot)_{L^{2}(\nu)}$. An increasing sequence of closed sets $\{F_{k}\}$ is called an $\mathscr{E}$-nest if $\cup_{k\geq 1}\mathscr{D}(F_{k})$ is $\mathscr{E}_{1}$-dense. A function $f$ is called $\mathscr{E}$-quasi-continuous if for any $\epsilon>0$, there exists an open set $\mathcal{O}$ with $\mathrm{Cap}(\mathcal{O})<\epsilon$ such that $f|_{S\backslash\mathcal{O}}$ is finite and continuous. We call a subset $\mathfrak{N}\subset S$ an $\mathscr{E}$-polar set if there exists an $\mathscr{E}$-nest $\{F_{k}\}$ such that $\mathfrak{N}\subset\cap_{k}(S\backslash F_{k})$.
@@
A Dirichlet form $(\mathscr{E},\mathscr{D})$ on $L^{2}(\nu)$ is quasi-regular if:
((Q1)) there exists an $\mathscr{E}$-nest $\{F_{k},k\geq 1\}$ consisting of compact sets;
((Q2)) there exists an $\mathscr{E}_{1}$-dense subset of $\mathscr{D}$ whose elements admit $\mathscr{E}$-quasi-continuous $\nu$-versions;
((Q3)) there exists $\{f_{k},k\geq 1\}\subset\mathscr{D}$ having $\mathscr{E}$-quasi-continuous $\nu$-versions $\{\tilde{f}_{k},k\geq 1\}\subset\mathscr{D}$ and an $\mathscr{E}$-polar set $\mathfrak{N}\subset S$ such that $\{\tilde{f}_{k},k\geq 1\}\subset\mathscr{D}$ separates the points on $S\backslash\mathfrak{N}$.